# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ 3.3 Biomass estimation
Additionally, a small number of studies use satellite LiDAR point clouds as weak supervision signals, combined with deep learning techniques, to map terrestrial biomass. Weber et al. [210] develop a deep learning model to predict global above-ground biomass density (AGBD), canopy height (CH), and canopy cover (CC) using 13-channel inputs from Sentinel-1/2, multispectral data, and DEMs. With GEDI LiDAR data as labels, the model employs a multi-head CNN architecture to simultaneously predict these variables and their uncertainties, enabling high-resolution global forest monitoring. Muszynski et al. [231] pre-train a Geospatial Foundation Model (GFM) with self-supervised learning, then fine-tune it on sparsely labeled data to estimate aboveground biomass (AGB) in Brazil. Using a Swin-B Transformer as the encoder, the model combines labeled and unlabeled data to improve generalization, achieving AGB estimation accuracy comparable to U-Net while reducing computational costs.
@@ 3.4 Summary and discussion
In remote sensing inversion tasks, traditional methods typically train models on small regions and apply them to the entire dataset. However, this approach often leads to significant errors when dealing with unseen data, affecting inversion accuracy. Using LiDAR data as weak supervision in combination with deep learning techniques and training jointly on large-scale datasets allows the model to learn broader features and shows great potential. However, existing research mainly focuses on network structure adjustments or pseudo-labeling, without fully leveraging weakly supervised learning techniques. Approaches such as contrastive learning, self-training, consistency loss, and active learning remain underexplored.
Moreover, remote sensing data are influenced by factors like geographic location, weather, seasonal variations, and sensor types, leading to significant domain shifts. Current research often assumes that data from different regions and times belong to the same domain, without fully utilizing domain adaptation techniques to address domain shift issues.
In conclusion, current remote sensing inversion methods largely rely on traditional statistical approaches, with relatively limited use of modern machine learning techniques, particularly weakly supervised learning, which can learn from imperfect data. Integrating weakly supervised learning into remote sensing inversion tasks can improve model accuracy and generalization and effectively address the complexities of remote sensing data, advancing inversion technologies.
@@ 4 Perspectives and Future Work
Leveraging Ubiquitous Annotations. The development of large-scale models, including generative models and vision-language foundation models, has enabled the generation of vast quantities of annotations [134]. These models have already demonstrated significant success in annotating remote sensing imagery. Beyond these automated approaches, numerous sources of annotations are readily available, including crowd-sourced data [233], network-generated labels [234], historical maps (e.g., OpenStreetMap [235]), and annotations derived from remote sensing images, as shown in Fig. 16. This abundance of annotation sources represents a valuable opportunity for advancing weakly supervised LiDAR interpretation. Future research should focus on fully harnessing these ubiquitous annotations to enhance the precision and efficiency of weakly supervised LiDAR-remote sensing tasks. By integrating these diverse annotation sources, researchers can reduce reliance on labor-intensive manual labeling, improve model performance, and significantly lower operational costs. The challenge lies in developing robust frameworks that can effectively utilize such diverse and sometimes noisy annotations to achieve consistent and scalable outcomes.
@@
Domain Variability in LiDAR Remote Sensing. The dynamic nature of these phenomena—driven by temporal, seasonal, and regional variability—exacerbates the challenges of domain shifts [16]. These challenges are further amplified by the inherent modality inconsistencies among LiDAR systems, such as airborne (ALS), terrestrial (TLS), and spaceborne sensors, which differ in resolution, coverage, and acquisition geometry. The combination of domain variability and multimodal inconsistencies makes obtaining comprehensive, consistent annotations both expensive and infeasible, particularly given the high cost and complexity of point cloud labeling.
@@
Weakly supervised learning offers a practical solution by leveraging limited labeled data and vast unlabeled datasets to mitigate annotation scarcity. Techniques such as semi-supervised learning [30, 31], feature alignment [23], and pseudo-labeling [32] are particularly suited to LiDAR data, utilizing its geometric and spatial properties. When combined with domain adaptation, these methods can address not only temporal and spatial domain shifts but also the modality gaps inherent to LiDAR systems.
@@
Future LiDAR research must prioritize annotation-efficient and modality-robust approaches to handle domain and sensor variability. Weak supervision has the potential to enhance model generalization and adaptability, enabling LiDAR to address dynamic Earth observation challenges such as disaster response, urban monitoring, and environmental change analysis. By bridging technical gaps, weakly supervised learning can ensure LiDAR’s broader applicability in real-world, rapidly evolving scenarios.
@@
LiDAR Remote Sensing Meets Open World. Current LiDAR remote sensing technologies have made significant progress in LiDAR interpretation and inversion. However, most existing methods focus on closed-set tasks, where neural networks are trained on a specific dataset and tested on the same category set. While effective for recognizing known classes, this approach struggles with the dynamic and uncertain nature of the real world, including unseen categories, environmental changes, and new tasks [236]. As a result, traditional closed-set methods are limited in addressing the diversity and complexity of the open world. Existing LiDAR datasets are relatively small and fixed in terms of categories. Future research should focus on developing LiDAR datasets with richer, more diverse categories that better reflect real-world scenarios. Additionally, integrating foundational models and multimodal data to develop cross-modal, generalizable, and scalable models will be a key direction for future work.
@@
New Paradigms for LiDAR Remote Sensing Retrieval. Most current approaches in LiDAR remote sensing involve extracting sparse signal-derived features, such as elevation or vegetation metrics, and using them to construct regression or classification models for large-scale remote sensing retrieval tasks. However, with the advancement of deep learning, there is growing interest in leveraging weak supervision to fuse LiDAR data with optical remote sensing imagery for improved retrieval accuracy [14]. Constructing spatiotemporally continuous remote sensing retrieval products using sparse supervision signals (e.g., LiDAR) [237]. This emerging paradigm raises two critical challenges: 1) how to effectively integrate multi-modal remote sensing data for retrieval tasks, and 2) how to utilize sparse LiDAR data as a weak supervisory signal to train deep learning models for large-scale retrieval.
@@
Meanwhile, the use of LiDAR as a weak supervisory signal introduces unique challenges in training deep learning models. Sparse LiDAR data, despite its limitations, provides valuable ground truth for guiding retrieval tasks across large scales. Research efforts should focus on designing frameworks that can generalize from sparse supervisory signals, incorporating techniques like semi-supervised learning, transfer learning, and domain adaptation. These methods have the potential to unlock new applications in areas such as vegetation biomass estimation, urban morphology reconstruction, and hydrological modeling.
@@
Datasets for LiDAR Remote Sensing. For large-scale environmental and geophysical parameter inversion, such as biomass estimation, LiDAR can serve as a critical weak supervision signal [210, 231]. However, the lack of comprehensive, large-scale LiDAR datasets tailored to these tasks limits its potential. High-resolution LiDAR data integrated with other remote sensing modalities could bridge this gap, enabling more accurate and scalable models for remote sensing applications [238].
@@
In the domain generalization context, the absence of datasets facilitating cross-modal, cross-season and cross-platform transfer learning poses significant challenges. Current datasets are predominantly developed for autonomous driving and lack the scope and diversity required for Earth observation. For example, cross-modal datasets enabling the transfer of knowledge between space-borne, airborne (ALS), and terrestrial (TLS) LiDAR remain unavailable. Such datasets are essential for studying domain gaps and advancing domain adaptation techniques for varying acquisition geometries and resolutions—key challenges in remote sensing.
Moreover, the scarcity of multimodal datasets combining LiDAR with optical, radar, or hyperspectral remote sensing data further constrains progress. Cross-modality datasets with aligned labels are critical for fostering weakly supervised learning and enabling models to leverage complementary information across sensing modalities.

### v2

@@ 4.3 Biomass estimation
In addition, a limited number of studies, under an incompletely supervised learning paradigm, employ satellite LiDAR point clouds as weak supervisory signals in combination with deep learning techniques for aboveground biomass (AGB) mapping. Weber et al. [233] develop a deep learning model to predict global AGB density, canopy height, and canopy cover using 13-channel inputs from Sentinel-1/2, multispectral data, and DEMs. With GEDI LiDAR data as labels, the model employs a multi-head CNN architecture to simultaneously predict these variables and their uncertainties, enabling high-resolution global forest monitoring. Muszynski et al. [264] pre-train a Geospatial Foundation Model (GFM) with self-supervised learning, then fine-tune it on sparsely labeled data to estimate AGB in Brazil. Using a Swin-B Transformer as the encoder, the model combines labeled and unlabeled data to improve generalization, achieving AGB estimation accuracy comparable to U-Net while reducing computational costs. Lyu et al. [265] address the sparsity issue of GEDI LiDAR data by innovatively employing a point-supervised U-Net model to convert continuous multi-source remote sensing imagery into a high-resolution, seamless forest aboveground carbon density map.
To enable large-scale, continuous mapping of forest aboveground biomass (AGB), recent studies primarily integrate multi-source remote sensing data with machine learning models. The core technical approach utilizes sparse footprint-level AGB density or canopy height data from NASA’s GEDI and ICESat-2 satellite LiDAR missions as weak supervisory signals, supplemented by diverse inputs such as Sentinel-2 or Landsat optical imagery, Sentinel-1 or ALOS-2 SAR backscatter, and SRTM-derived terrain variables, to achieve seamless mapping through machine learning. Among modeling approaches, Random Forest is most commonly employed due to its robustness in handling high-dimensional and heterogeneous data, while alternative models include Gradient Boosting Machines (GBM), Support Vector Regression (SVR), and convolutional neural networks (CNNs) such as U-Net, which more effectively capture spatial context. Despite significant progress, several key challenges remain, including signal saturation in high-biomass forests leading to underestimation, geolocation and spatial mismatches between sparse LiDAR footprints and high-resolution imagery, limited model transferability across ecological regions, and the complexity of multi-sensor data fusion caused by differences in spatial and temporal resolutions.
Figure 16: Regional bathymetric map of Shanhu Island and Yuanzhi Island. The image is obtained from [266].
4.4 Water depth estimation
Traditional bathymetry relies on in-situ methods such as shipborne sonar, which are hindered by low efficiency and high labor and time costs, limiting their suitability for large-scale, high-frequency monitoring. Remote sensing inversion techniques provide a rapid, cost-effective alternative for obtaining underwater terrain data, thus garnering broad attention. In particular, ICESat-2’s unprecedented high resolution and elevation accuracy enable revolutionary remote sensing bathymetry in the absence of field measurements, offering extensive application potential, as shown in Fig. 16.
Existing studies predominantly employ domain generalization approaches. The dominant paradigm entails training machine learning or deep learning models using discrete LiDAR-derived labels together with continuous remote sensing imagery, and subsequently applying the trained models to large-scale bathymetric mapping. Thomas et al. [267] employ an extremely randomized tree regression model to learn the relationship between precise bathymetric points (labels) from ICESat-2 LiDAR and multispectral features from Landsat-8 imagery, thereby generating high-precision continuous bathymetric maps. To enhance accuracy, Peng et al. [268] propose a physics-assisted convolutional neural network (PACNN) that embeds water radiative transfer principles into the network architecture and uses precise ICESat-2 depth points as training labels for Sentinel-2 multispectral imagery, substantially improving the precision of traditional remote sensing bathymetric mapping. Yang et al. [269] innovatively integrate ICESat-2 along-track depth profiles with spectral trends from Sentinel-2 imagery to construct virtual “free profiles” that augment training data, then apply Random Forest and other machine learning inversion models to achieve high-precision lake-wide bathymetric mapping without in situ measurements. Similarly, Othman et al. [270] combine Sentinel-2 multispectral imagery and ICESat-2 LiDAR data with AI methods such as quantile regression forests (QRF) to estimate shallow-water depths in Darbandikhan Lake, Iraq, without relying on conventional ground surveys. For global lakes, Lv et al. [271] leverage piecewise gradient boosting (piecewise GB) integrated with satellite imagery and LiDAR to deliver dynamic, continuous estimates of water depth and volume, particularly for small- to medium-sized lakes. Gülher et al. [272] use multi-source inputs comprising optical satellite imagery, ICESat-2 LiDAR, and conventional echo-sounder data, along with Random Forest and XGBoost models, to overcome limitations in traditional shallow-water mapping and produce seabed bathymetry maps meeting IHO CATZOC A1 standards. Du et al. [273] train a Kolmogorov–Arnold network (KAN) using precise ICESat-2 depth points and apply it to Sentinel-2 imagery to retrieve bathymetric maps comparable to those from in situ surveys. Zhao et al. [266] exploit the broad coverage of multispectral imagery and the sparse, high-precision measurements from spaceborne LiDAR, treating LiDAR data as exact labels to train an error-correction model that refines large-scale initial inversions and yields precise regional underwater topography. Xie et al. [274] develop a physics-informed convolutional neural network (PI-CNN) that incorporates radiative transfer principles, using ICESat-2 depth points as labels and wide-coverage Sentinel-2 multispectral imagery as input features to map nearshore water depths over extensive areas. Zhang et al. [275] utilize a Bayesian-optimized CatBoost model, training it with precise ICESat-2 depth data as labels to interpret Sentinel-2 imagery and address challenges in large-scale, automated, high-precision bathymetry for supraglacial lakes. Liu et al. [276] first perform seabed classification to decouple mixed signals from water and seabed, then apply a neural network model fusing ICESat-2 LiDAR data with Sentinel-2 optical imagery to tackle bathymetric inversion challenges in complex shallow-water environments.
The integration of spaceborne LiDAR with optical imagery has become the prevailing paradigm for bathymetric inversion. This approach leverages high-precision, along-track sparse depth measurements from spaceborne LiDAR (e.g., ICESat-2) as reference labels together with spectral features derived from high spatiotemporal–resolution optical imagery (e.g., Sentinel-2) to train machine-learning models (e.g., Random Forest, convolutional neural networks) for regional, continuous depth estimation. Despite its considerable promise, the method faces several critical challenges in practice: (1) Poor model transferability. Models calibrated for a given region are strongly dependent on the local inherent optical properties of the water and on substrate composition. When transferred to areas with different optical regimes, model performance commonly deteriorates, limiting generalizability. (2) Spatial resolution mismatch among data sources. Scale discrepancies between the LiDAR laser footprint ($\approx$17 m for ICESat-2) and the pixel size of typical optical sensors (e.g., 10 m for Sentinel-2) introduce ambiguity when assigning sparse LiDAR depth labels to dense image pixels, thereby increasing label uncertainty. (3) Data sparsity and physical limitations. The along-track sampling pattern of ICESat-2 produces incomplete spatial coverage. Furthermore, both active and passive optical signals attenuate rapidly in deep (>20–30 m) or highly turbid waters, constraining effective penetration depths and producing systematic underestimation of depths in deeper areas. Environmental factors such as sunglint and atmospheric contamination further degrade data quality. (4) Difficulties in separating water-column and substrate contributions. Satellite-measured signals comprise a composite of water column backscattering and bottom reflectance. In environments with heterogeneous substrates (e.g., mixtures of sand, coral, and seagrass), reliably disentangling these components remains challenging and substantially affects inversion accuracy.
4.5 Other applications
Canopy cover mapping. Canopy cover, defined as the proportion of ground obscured by the vertical projection of tree crowns, serves as a core indicator for assessing forest health, modulating ecosystem functions, and evaluating habitat quality. Existing studies primarily employ generalization approaches, with the prevailing paradigm involving the training of machine learning or deep learning models using discrete LiDAR data as labels alongside continuous remote sensing imagery, followed by the application of these models to large-scale mapping. Narine et al. [277] compare ICESat-2 satellite data against high-precision airborne LiDAR to assess two canopy cover estimation methods—direct computation and Random Forest—highlighting the superior performance of machine learning. Viana-Soto et al. [278] train a support vector regression (SVR) model with airborne LiDAR-derived labels for forest structure parameters (e.g., canopy cover and height), incorporating long-term Landsat imagery features to reconstruct historical metrics—including vegetation cover (VC), tree cover (TC), mean height (MH), and height variation coefficient (CVH)—spanning the past 30 years. Similarly, Schlickmann et al. [279] fuse multi-source data by using GEDI data and Sentinel-1 SAR and Harmonized Landsat-Sentinel-2 (HLS) optical imagery to train a Random Forest model, yielding the first high-resolution, seamless statewide canopy cover map for Florida.
Carbon stock estimation. Forest carbon stock represents the total carbon absorbed and stored by forest ecosystems, playing a crucial role in mitigating global climate change. Traditional plot surveys calculate individual tree carbon stocks through field measurements of tree attributes combined with allometric equations, offering high accuracy but incurring substantial costs and scalability challenges. Remote sensing methods, by contrast, leverage satellite imagery—such as optical or radar data—to construct models that estimate regional-scale carbon stock distributions, affording extensive spatial coverage. Existing studies primarily employ generalization approaches, with the prevailing paradigm involving the training of machine learning or deep learning models using discrete LiDAR data as labels alongside continuous remote sensing imagery, followed by the application of these models to large-scale mapping. Jiao et al. [280] fuse Sentinel-2 optical imagery with sparse GEDI data using a fully convolutional network (FCN) to generate high-resolution continuous forest canopy height maps. They then apply an airborne LiDAR (ALS)-derived canopy height–carbon stock model to convert these maps into high-precision carbon stock distributions for the project area, from which they estimate forest carbon emission reductions. Sun et al. [281] integrate limited field plots with high-precision ALS data through a Bayesian hierarchical model to produce reliable carbon stock samples in representative areas; these serve as high-quality labels to train a Random Forest model, facilitating high-precision mapping of forest carbon stocks across larger regions using broad-coverage Sentinel-2 and auxiliary data.
In addition, Araza et al. [282] employ an ensemble machine learning model (integrating Random Forest, gradient boosting trees, and support vector machines) to learn relationships between multi-source satellite remote sensing data (e.g., global biomass and forest height change maps) and precise carbon flux labels from national forest inventories and airborne LiDAR, thereby generating continuous carbon flux maps for five countries spanning 2010–2018 to support United Nations carbon accounting. Xi et al. [283] leverage phenological differences between overstory evergreen conifers and understory deciduous vegetation, training a support vector regression model with multi-temporal Sentinel-2 optical imagery and GEDI LiDAR data to enable large-scale mapping of spatially continuous understory vegetation density. Similarly, Yu et al. [284] propose a large-scale method for estimating forest growing stock volume, using high-precision airborne LiDAR (ALS) data as a “bridge” to generate reliable training labels for Sentinel-2 imagery and producing high-accuracy maps of standing timber volume for larch forests across China via a Random Forest model. Yao et al. [285] train a Random Forest model to capture complex relationships between ICESat-2 satellite elevation points and surface features from Sentinel-1/Sentinel-2 imagery, generating continuous high-precision terrain elevation maps at 10 m resolution.
These studies typically employ a domain generalization strategy, where models are first trained on small-scale regions and subsequently generalized to unseen large-scale datasets. They leverage spatially sparse yet high-precision, weakly supervised signals from spaceborne LiDAR (e.g., GEDI, ICESat-2) footprints to provide accurate measurements of vertical forest structure, such as canopy height and elevation, despite their discontinuous coverage. Concurrently, ALS data often serve as a bridge to upscale local ground-truth plot data, generating denser pseudo-labels for model training. In terms of model selection, Random Forest is widely adopted due to its effectiveness in handling high-dimensional data, its robustness against overfitting, and its stable performance. Overall, the primary research challenges include the signal saturation of optical remote sensing (which makes it difficult to distinguish subtle biomass differences in dense vegetation), the spatiotemporal mismatch of multi-source data (such as alignment errors between LiDAR footprints and satellite pixels), and the insufficient representativeness of reference data. These factors introduce significant uncertainty when extrapolating results to unsampled regions, highlighting fundamental challenges in domain adaptation and generalization for large-scale LiDAR–remote sensing.
@@ 4.6 Summary and discussion
LiDAR-based inversion seeks to generalize from sparse spatiotemporal observations to continuous, large-scale parameter maps—a problem that is essentially domain generalization (DG): models are learned from limited measurements in observed regions or periods and applied to unseen domains. supervisory signals typically come from two hierarchical sources: (i) in situ ground measurements (e.g., tree height, biomass, canopy cover), which provide the most accurate references; and (ii) LiDAR data (e.g., GEDI, ALS), which supplies structurally informative signals bridging sparse measurements and continuous optical imagery. These complementary properties make LiDAR an effective weak supervision source for large-scale inversion of biomass, tree height, water depth, canopy cover, and carbon stock.
However, current inversion pipelines face coupled data–model limitations. On the data side, LiDAR samples are spatially incomplete and unevenly distributed, subject to geolocation errors and canopy-structure–dependent biases; RH metrics are statistical proxies rather than physical ground truth, and ICESat-2 exhibits increased sensitivity to terrain and noise. A frequently overlooked issue is resolution mismatch: for instance, a single GEDI footprint ( 25 m diameter) is often assigned to the center pixel of a 10 m Sentinel-2 image (or finer), implicitly assuming within-footprint homogeneity and ignoring sub-footprint heterogeneity. Similar mismatches between airborne LiDAR point clouds and imagery introduce systematic supervision noise and propagate uncertainty through training.
On the modeling/method side, many large-scale studies still rely on simple statistical or classical ML models (e.g., linear regression, Random Forest, Gradient Boosting) that are easy to train under sparse labels but offer limited cross-domain generalization. Although CNNs and CNN–Transformer hybrids are increasingly used to exploit spatial context (e.g., Landsat/Sentinel plus DEM), label uncertainty and domain shift are often not treated explicitly. In practice, model transferability is typically enhanced through heuristic and experience-based strategies: incorporating empirical biogeophysical variables (e.g., NDVI, topography, tidal dynamics), region-specific training or local fine-tuning, temporal stratification to reduce seasonal domain shifts, and multi-source data fusion (e.g., combining ICESat-2, optical imagery, SAR, and tidal information). Temporal sample migration is also used to enlarge training sets and improve robustness across environmental conditions. While these strategies can enhance performance within local regions, they depend heavily on expert knowledge, manual rules, and task-specific tuning, which limits scalability and reproducibility. As a result, most existing applications achieve only regional or site-specific generalization, and their accuracy tends to degrade substantially when extended to broader spatial or temporal domains.
In contrast, advances in machine learning offer systematic, learnable DG solutions—such as adversarial domain alignment, domain-invariant feature learning, test-time adaptation, meta-learning, and uncertainty-aware representation learning—that can replace or augment these heuristic strategies. Weak supervision provides the algorithmic foundation for treating LiDAR observations as imperfect signals rather than perfect labels: pseudo-label generation, consistency regularization, and contrastive learning improve robustness to sparse and noisy supervision; multi-modal fusion and self-training support scale transfer; active learning can prioritize informative LiDAR samples. Combining weak supervision with DG/DA/TTA enables models to adapt to spatiotemporal heterogeneity, resolution mismatch, and label uncertainty in a principled and scalable manner. This shift from heuristic rules to learnable generalization represents a critical next step toward robust, transferable, and cost-efficient LiDAR-based inversion.
@@ 5 Perspectives and Future Work
5.1 Challenges of applying WSL to LiDAR remote sensing
Uniqueness of LiDAR remote sensing data. Compared to natural images and remote sensing imagery, LiDAR remote sensing differs fundamentally not only in data organization (raster grids vs. point clouds/waveforms) but also in exhibiting more complex and challenging characteristics. First, point cloud sampling is influenced by multiple factors such as altitude, scanning angle, echo energy, and terrain occlusion, resulting in highly uneven density with sparse, irregular, and spatially discontinuous features; this contrasts sharply with the fixed resolution of imagery and the high density and local stability of indoor point clouds. Second, the wide range of observation platforms (e.g., drones) leads to substantial variations in viewpoint and scale, causing much greater heterogeneity in observation geometry and object structures than in imagery or indoor scenes. Third, annotation is significantly more difficult due to diverse object types, complex terrain undulations, and occlusions, whereas indoor point clouds benefit from simple structures and controllable scenes, facilitating standardized and automated labeling. Finally, LiDAR domain variations are strongly coupled with geographic environments, seasons, and ecological dynamics, with temporal changes markedly amplified; in contrast, domain differences in natural images primarily stem from illumination styles, remote sensing imagery is governed by imaging geometry, and indoor point clouds exhibit the most stable environments.
In summary, the challenges of LiDAR remote sensing in sampling sparsity, geometric scale variations, annotation complexity, and spatiotemporal domain heterogeneity far exceed those of images and indoor point clouds, directly increasing the difficulty of algorithm modeling.
Limitations of existing LiDAR WSL. These characteristics dictate that weakly supervised learning for LiDAR remote sensing cannot simply adopt techniques from images or indoor point clouds. First, the discrete and sparse nature of point clouds disrupts the spatial continuity assumptions underlying image-based weak supervision, making label propagation and pseudo-label generation unstable in discontinuous regions, whereas indoor point clouds are less affected due to their density and controllability. Second, LiDAR lacks spectral and textural priors, with dominant information being geometric structures and intensity, rendering image-based weak supervision methods reliant on saliency or attention difficult to transfer; indoor point clouds often incorporate RGB data, enabling direct leverage of visual foundation models. Third, its domain variations and spatiotemporal dynamics are far stronger than in images and indoor point clouds, imposing higher demands on model robustness and generalization. LiDAR has a more urgent need for domain generalization (DG), yet research in this direction remains scarce; in contrast, the image domain has mature methods and pre-trained models (e.g., CLIP, SAM) to mitigate domain shifts, while LiDAR lacks comparable support. Finally, LiDAR’s unique waveform-point cloud multimodal features (e.g., waveform decomposition, multiple echoes, sub-footprint) further increase uncertainty in data interpretation and pseudo-label modeling.
Therefore, the weak supervision problem in LiDAR remote sensing is not merely “insufficient labeling” but a systemic challenge driven by sampling structures, feature modalities, spatiotemporal dynamics, and task mechanisms, necessitating the development of a dedicated weak supervision theoretical and methodological framework independent of images and indoor point clouds.
5.2 LiDAR remote sensing meets foundation models
Foundation models, characterized by large-scale, multi-source pre-training, are driving a paradigm shift in artificial intelligence. These models, including Large Language Models (LLMs) and Vision-Language Models (VLMs), demonstrate remarkable generalization and zero/few-shot learning capabilities, enabling adaptation to a wide range of downstream tasks through fine-tuning. LLMs excel at the deep understanding and generation of natural language [287], while VLMs establish profound connections between visual content (images, videos) and textual descriptions by aligning visual and textual modalities.
In the field of Earth observation, the deep integration of LiDAR technology with foundation models is forging a new path for 3D perception. This convergence combines the precise 3D geometric measurement capabilities of LiDAR with the powerful deep semantic reasoning of foundation models, advancing 3D scene understanding from traditional classification of predefined categories toward an open-world cognitive intelligence paradigm that supports natural language interaction. This fusion exhibits significant advantages in two core areas: Parameter Inversion: In tasks such as forest biomass estimation, high-precision geometric priors from LiDAR (e.g., canopy structure) effectively constrain numerical models. When combined with the semantic context and domain knowledge extracted by VLMs/LLMs, the accuracy, robustness, and generalization of model estimations across different regions and sensors are significantly enhanced [288]. Scene Interpretation: For tasks like semantic segmentation and object detection, the open-vocabulary characteristic of VLMs overcomes the reliance of traditional models on fixed label sets. This enables open-set recognition and rare object detection, dramatically increasing the flexibility and intelligence of scene interpretation [289].
However, transferring the powerful semantic capabilities of foundation models to sparse, irregular LiDAR 3D point cloud data remains a formidable challenge. Mainstream foundation models are predominantly trained on 2D regular image grids and 1D text sequences, rendering their network architectures incompatible with direct point cloud processing. This incongruity leads to two primary issues: Modality Mismatch and Projection Error: The significant discrepancy between the 2D image-based pre-training of foundation models and the sparse 3D geometry of LiDAR point clouds results in performance degradation. Furthermore, projecting point clouds onto 2D planes (e.g., elevation maps, bird’s-eye views) inevitably introduces geometric information loss, occlusion distortion, and resolution limitations, which further compromises the accuracy of downstream tasks. Cross-Modal Alignment Uncertainty and Domain Shift: When transferring knowledge using registered image-LiDAR data pairs [290], noise and semantic ambiguity from weak labels, coupled with unreliable alignment due to temporal, viewpoint, resolution, and geometric registration discrepancies, diminish the effectiveness of knowledge transfer. Moreover, significant domain shifts, arising from sensor heterogeneity and geographic variability, severely constrain the generalization performance of these models.
To systematically bridge the modal gap between LiDAR and foundation models, the academic community is actively exploring three emerging strategies: (i) Prompt/Adapter-based Fine-tuning: This approach involves lightweight adjustments to inject structural cues extracted from LiDAR (e.g., canopy height profiles) as geometric prompts into pre-trained models. (ii) Cross-modal Pre-training: This strategy aligns the geometric features of LiDAR with the semantic features of imagery within a shared representation space during the pre-training phase, using methods like contrastive learning or knowledge distillation. (iii) Pseudo-label Refinement: This iterative process fuses coarse semantic labels generated by VLMs with precise geometric information from LiDAR to produce high-quality supervisory signals that, in turn, refine the model’s training.
Within this new paradigm, the role of weak supervision extends beyond merely addressing label scarcity; it is elevated to a critical bridge connecting LiDAR’s geometric information with the semantic space of foundation models. Weak supervisory signals derived from LiDAR (e.g., canopy height, digital elevation models) serve as a low-cost source of rich structural information. They not only anchor the cross-modal feature alignment but also inject essential geometric priors into the high-dimensional semantic space. This guides large models to comprehend 3D geometric structures without requiring massive volumes of pre-trained point cloud data. For instance, AlphaEarth Foundations utilizes sparse GEDI LiDAR data (for forest canopy height) as a reconstruction target within its self-supervised learning framework. This process enables the model to leverage other, denser data sources, such as optical and radar imagery, to learn and generate continuous, high-resolution, global estimates of vegetation height [291]. In return, foundation models endow LiDAR analysis tasks with unprecedented semantic breadth and open-world reasoning capabilities. This synergistic cycle of geometric priors and semantic capabilities establishes a bidirectional pathway for knowledge transfer, laying a solid foundation for the ultimate development of point cloud-native foundation models.
5.3 Future directions for LiDAR remote sensing
@@
Leveraging Ubiquitous Annotations. The development of large-scale models, including generative models and vision-language foundation models, has enabled the generation of vast quantities of annotations [172]. These models have already demonstrated significant success in annotating remote sensing imagery. Beyond these automated approaches, numerous sources of annotations are readily available, including crowd-sourced data [292], network-generated labels [293], historical maps (e.g., OpenStreetMap [294]), and annotations derived from remote sensing images, as shown in Fig. 17. This abundance of annotation sources represents a valuable opportunity for advancing weakly supervised LiDAR data interpretation. Future research should focus on fully harnessing these ubiquitous annotations to enhance the precision and efficiency of weakly supervised LiDAR-remote sensing tasks. By integrating these diverse annotation sources, researchers can reduce reliance on labor-intensive manual labeling, improve model performance, and significantly lower operational costs. The challenge lies in developing robust frameworks that can effectively utilize such diverse and sometimes noisy annotations to achieve consistent and scalable outcomes.
@@
Domain Variability in LiDAR Remote Sensing. The dynamic nature of these phenomena—driven by temporal, seasonal, and regional variability—exacerbates the challenges of domain shifts [18]. These challenges are further amplified by the inherent modality inconsistencies among LiDAR systems, such as ALS, TLS, and spaceborne sensors, which differ in resolution, coverage, and acquisition geometry. The combination of domain variability and multimodal inconsistencies makes obtaining comprehensive, consistent annotations both expensive and infeasible, particularly given the high cost and complexity of point cloud labeling.
@@
WSL offers a practical solution by leveraging limited labeled data and vast unlabeled datasets to mitigate annotation scarcity. Techniques such as semi-supervised learning [55, 56], feature alignment [92], and pseudo-labeling [6] are particularly suited to LiDAR data, utilizing its geometric and spatial properties. When combined with DA, these methods can address not only temporal and spatial domain shifts but also the modality gaps inherent to LiDAR systems.
@@
Future LiDAR research must prioritize annotation-efficient and modality-robust approaches to handle domain and sensor variability. Weak supervision has the potential to enhance model generalization and adaptability, enabling LiDAR to address dynamic Earth observation challenges such as disaster response, urban monitoring, and environmental change analysis. By bridging technical gaps, WSL can ensure LiDAR’s broader applicability in real-world, rapidly evolving scenarios.
@@
LiDAR Remote Sensing Meets Open World. Current LiDAR remote sensing technologies have made significant progress in LiDAR data interpretation and LiDAR-based inversion. However, most existing methods focus on closed-set tasks, where neural networks are trained on a specific dataset and tested on the same category set. While effective for recognizing known classes, this approach struggles with the dynamic and uncertain nature of the real world, including unseen categories, environmental changes, and new tasks [295]. As a result, traditional closed-set methods are limited in addressing the diversity and complexity of the open world. Existing LiDAR datasets are relatively small and fixed in terms of categories. Future research should focus on developing LiDAR datasets with richer, more diverse categories that better reflect real-world scenarios. Additionally, integrating foundational models and multimodal data to develop cross-modal, generalizable, and scalable models will be a key direction for future work.
@@
New Paradigms for LiDAR Remote Sensing Inversion. Most current approaches in LiDAR remote sensing involve extracting sparse signal-derived features, such as elevation or vegetation metrics, and using them to construct regression or classification models for large-scale remote sensing retrieval tasks. However, with the advancement of deep learning, there is growing interest in leveraging weak supervision to fuse LiDAR data with optical remote sensing imagery for improved retrieval accuracy [16]. Constructing spatiotemporally continuous remote sensing retrieval products using sparse supervisory signals (e.g., LiDAR) has emerged as a promising research direction [296]. This emerging paradigm raises two critical challenges: 1) how to effectively integrate multi-modal remote sensing data for retrieval tasks, and 2) how to utilize sparse LiDAR data as a weak supervisory signal to train deep learning models for large-scale retrieval.
@@
Meanwhile, the use of LiDAR as a weak supervisory signal introduces unique challenges in training deep learning models. Sparse LiDAR data, despite its limitations, provides valuable ground truth for guiding retrieval tasks across large scales. Research efforts should focus on designing frameworks that can generalize from sparse supervisory signals, incorporating techniques like semi-supervised learning, transfer learning, and DG. These methods have the potential to unlock new applications in areas such as vegetation biomass estimation, urban morphology reconstruction, and hydrological modeling.
@@
Datasets for LiDAR Remote Sensing. For large-scale environmental and geophysical parameter inversion, such as biomass estimation, LiDAR can serve as a critical weak supervisory signal [233, 264]. However, the lack of comprehensive, large-scale LiDAR datasets tailored to these tasks limits its potential. High-resolution LiDAR data integrated with other remote sensing modalities could bridge this gap, enabling more accurate and scalable models for remote sensing applications  [297].
@@
In the domain generalization context, the absence of datasets facilitating cross-modal, cross-season and cross-platform transfer learning poses significant challenges. Current datasets are predominantly developed for autonomous driving and lack the scope and diversity required for Earth observation. For example, cross-modal datasets enabling the transfer of knowledge between space-borne, ALS, and TLS remain unavailable. Such datasets are essential for studying domain gaps and advancing DA techniques for varying acquisition geometries and resolutions—key challenges in remote sensing. Moreover, the scarcity of multimodal datasets combining LiDAR with optical, radar, or hyperspectral remote sensing data further constrains progress. Cross-modality datasets with aligned labels are critical for fostering WSL and enabling models to leverage complementary information across sensing modalities.